# The human-machine paradox:

### v5

@@ Abstract
When deploying artificial skills, managers widely assume that combining them with the human factor is a safe harbor, mitigating the risks of full automation in high-complexity tasks. This paper formally challenges the economic validity of this widespread assumption, arguing that the true bottom-line economic utility of a human-machine skill policy is dangerously misunderstood and highly contingent on situational and design factors. To investigate this gap, we develop an in-silico framework based on Monte Carlo simulations grounded in empirical pragmatism to quantify the economic impact of human and machine skills in the execution of tasks presenting varying levels of complexity. Our results show that a human-machine strategy can yield the highest economic utility in complex scenarios, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine approach can perform worse than either the machine-exclusive or the human-exclusive policy, actively destroying value under the pressure of costs that are not compensated by sufficient performance gains. The takeaway for decision-makers is unambiguous: when the context is complex and critical, simply allocating human and machine skills to a task may be insufficient, and far from being a silver-bullet solution or a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of skill policy decisions ex ante. This provides a foundation for scaling experiments and study more general policy implications as candidates for future empirical validation.
@@ 1 Introduction
Prominent research analyzing the impact of technology over skill demand shows that computer capital allocations tend to result in automating well-defined and rule-based routine tasks, complementing human skills in tasks requiring abstraction, problem-solving, or creativity (Autor et al., 2003), and supplementing human judgment (Autor, 2024), with experimental evidence in (Peng et al., 2023), (Noy and Zhang, 2023), and (Brynjolfsson et al., 2025), provided the ability to use artificial capabilities effectively (Agarwal et al., 2023) (Autor, 2024). Meanwhile, technological advancements in methods, combined with increasingly affordable computational capabilities, have allowed artificial skills to achieve remarkable performance in tasks where automation has historically struggled. Notable examples include image recognition (Krizhevsky et al., 2012), natural language processing (Vaswani et al., 2017), game play (Silver et al., 2017), skin cancer diagnostics (Esteva et al., 2017), recommendation systems (He et al., 2017) and protein-folding research (Jumper et al., 2021), just to mention some widely recognized examples. Moreover, language models deployed through deep neural networks (Brown et al., 2020) (Hoffmann et al., 2022) (Team et al., 2023) (Touvron et al., 2023) have contributed to amplify expectations regarding the present and future performance of artificial skills in the execution of a wide variety of cognitive tasks, making them extremely accessible to a broad audience through user-friendly interfaces. These examples emphasize the potential of technological progress to expand the set of tasks that can be automated (Acemoglu, 2025). However, when tasks are characterized by high levels of generalization difficulty Chollet (2019), machines can struggle to achieve high levels of performance, presumably due lack of the abstraction and analogy-making abilities typical of the human mind. Nevertheless, while human skills may outperform artificial skills when facing surprising task content, they may be less efficient and less scalable. Then, an alternative solution to the human-machine dichotomy is to combine these two skills to get the best of both, enhancing execution performance at a cost of the dual skill structure. From an organizational perspective, these conditions make the administrative behavior (Simon, 1947) required to determine optimal capital and labor allocations far from straightforward. We call these skill policy decisions, which are particularly interesting in case of cognitive tasks, because compartmentalizing them is harder than manual or routine ones. To analyze this problem, we combine a task-based framework from the theory of Economics with provisions from Computer Science and Statistics into a model suitable to generate synthetic data through Monte Carlo simulations. Then, we leverage synthetic data to analyze the economic effects of skill policy decisions, from task-execution performance to bottom-line utility, in relation with initial conditions. To this end, we assume that task performance is not a deterministic outcome, but an emergent property shaped by complex interactions between skills and tasks under varying levels of complexity. Our goal is to understand more about the factors that can influence the economic effects of skill policy decisions, by leveraging bottom-up simulation to quantify emerging insights that may be transparent to top-down analysis. In this context, the computer science literature has extensively documented the technical challenges of generalization and the performance limitations of models on out-of-distribution or complex tasks. Therefore, our contribution is not to simply rediscover this qualitative insight. We introduce a formal in-silico framework that translates these known technical performance characteristics into quantifiable economic utility measures. To our knowledge, no prior work provides a formal model for ex-ante evaluation of skill policies that explicitly bridges the technical concept of generalization difficulty with the economic concepts of error costs and bottom-line economic utility. Thus we genuinely believe that our proposed framework can be a novel and useful tool for decision-makers, expanding the discussion from skill performance alone to skill policy, and to how the skill policy of choice should be deployed to maximize economic utility.
2 Referential frameworks
Our work builds upon a task-based model of production (Farboodi and Veldkamp, 2021) characterizing the quality-adjusted output of a firm $i$ at time $t$ as $y_{i,t}=A_{i,t}k_{i,t}^{\alpha}$, where $k_{i,t}^{\alpha}$ represents the units of output of quality $A_{i,t}$ produced by leveraging $k_{i,t}$ units of capital, and $A_{i,t}=g\big((a_{i,t}-a_{i,t}^{opt})^{2}\big)$ is function of the squared distance between $a_{i,t}$ (the production technique of choice for firm $i$ at time $t$) and the optimal technique $a_{i,t}^{opt}=\theta_{t}+\epsilon_{a,i,t}$}, with $\theta_{t}$ and $\epsilon_{a,i,t}$ being the jointly-observed persistent and transitionary components of the optimal technique, and $g$ a strictly decreasing function. This formulation, by means of the introduction of the squared distance between the actual and the optimal production skill, makes the role of technological performance explicit in the determination of production output of automated tasks. Also, this framework complements the treatments illustrated in (Autor and Thompson, 2025), where a worker is able to execute a task only if the worker’s expertise level is equal or higher than the one required by the task (the hierarchical expertise assumption), and automation deploys below a certain threshold of task complexity (the hierarchical automation assumption), a condition also illustrated in (Acemoglu, 2025) as effect of the comparative advantage offered by labor with respect to capital in newer or complex tasks. Interestingly, this task-based model aligns with the idea of generalization difficulty as presented in (Chollet, 2019), i.e. the increased complexity introduced by additional levels of uncertainty presenting at execution time with respect to learning time. The compatibility of these seminal models constitutes the foundational ground for our work.
@@ 3 Model
By combining the provisions of the referential frameworks introduced in Section 2, we conceptualize performance as a stochastic property that emerges from the interaction between task content and skills, conditioned upon local and global generalization difficulty, extending the definition of production technique to include models operated by the human mind. In this context, we focus on cognitive tasks, leaving the opportunity to relax this assumption in further steps. Herein, we move by abstraction to defining the distance between the actual and optimal skills in terms of loss ($\mathcal{L}$), such that the quality $A_{i,t}$ can be expressed as $A_{i,t}=g\big(\mathcal{L}_{i,t}\big)$, where $\mathcal{L}_{i,t}=f(a_{i,t},a_{i,t}^{opt})$ and $f$ is any appropriate function measuring the absolute gap between the actual and optimal technological skills. By doing this, we open the possibility to consider alternative measures of distance, including divergence or dissimilarity (also, if necessary, multiple measures of loss may be combined). Then, when the range of potential values of $\mathcal{L}_{i,t}$ is known, we can obtain a normalized loss $\mathcal{L}^{\prime}_{i,t}\in[0,1]$ through the normalization process of choice (for example, $\mathcal{L}^{\prime}_{i,t}=[\mathcal{L}_{i,t}-\min(\mathcal{L}_{i,t})]/[\max(\mathcal{L}_{i,t})-\min(\mathcal{L}_{i,t})]$), and a normalized performance $\theta_{i,t}=(1-\mathcal{L}^{\prime}_{i,t})$. This allows to formally transition from the measure of technological distance to the measure of performance, with consequently easier operational deployment. Based on this initial framing, we implement a computational pipeline that, given initial conditions at time $t$, leverages a stochastic process to determine the normalized performance $\theta_{i,c,d,e}$ obtained by the firm $i$ deploying the skill policy $c$ in the execution of a task with difficulty $d$ at time $t+e$, where $e\in[0,E-1]$ is a task-execution epoch. To this end, without sacrificing generality, our design implements three skill policies and three levels of task difficulty. The skill policy options are the human-exclusive ($H$, relying on labor allocations only), the machine-exclusive ($M$, i.e. full automation), and the human-machine ($HM$, combining human and machine skills). The implementation of local generalization difficulty is probabilistic, with distributions $P(\theta)_{c,d,e}$ conditioned on $c$, $d$, and $e$. More specifically, the local generalization difficulty is inherently described by each individual probability distribution $P(\theta)_{c,d,e}$, while the global generalization difficulty is modeled by having different distributions for different levels of difficulty $d$. At each epoch, the performance obtained by the skill policies $H$ or $M$ (denoted as $\theta_{i,c=H,d,e}$ and $\theta_{i,c=M,d,e}$, and more synthetically as $\theta_{H}$ and $\theta_{M}$, respectively) is randomly drawn according to the appropriate probability distribution $P(\theta)_{c,d,e}$. When, instead, the skill policy adopted by the firm is $HM$, the performance computes as $\theta_{HM}=a(\theta_{H},\theta_{M},\gamma_{HM})$, where $a$ is an arbitrary function defining how the individual skills $H$ and $M$ interact, depending also on an augmenting factor $\gamma_{HM}$. Herein, the concept of augmentation evokes amplification (Ashby, 1956), collaboration (Licklider, 1960) and integration (Engelbart, 1962), characterizing a human-machine skill policy in ways that go from task decomposition and sub-task delegation to complex collaborative partnerships. Then, based on the performance $\theta_{i,c,d,e}$, the quality-adjusted production output computes as $y_{i,c,d,e}=g^{\prime}\big(\theta_{i,c,d,e}\big)$, where $g^{\prime}:\theta_{i,c,d,e}\in[0,1]\rightarrow y_{i,c,d,e}\in[0,1]$ is a monotone and non decreasing function (the combined evaluation of $A_{i,t}$ and ${kl}^{\alpha}_{i,t}$ through the function $g^{\prime}$ is a desired behavior of our model, to streamline the formal relationship between task-execution performance and quality-adjusted output). Next, the value of the production output is determined as $v_{i,c,d,e}=max(0,min(1,y_{i,c,d,e}*mc_{c}*\delta_{c,e}))$, where $mc_{c}\in[0,1],\forall c$ is the margin of contribution associated to each skill policy, with point values $mc_{H},mc_{M},mc_{HM}$, and $\delta_{c,e}$, which depends from both $c$ and $e$, represents the skill-specific evolution of the margin of contribution over epochs. To this end, $\delta_{c,e}$ is computed by linear interpolation as $\delta_{c,e}=interpolate(1,1+\delta_{c},e,E)$, according to Algorithm 3, where $\delta_{c}=\{\delta_{H},\delta_{HM},\delta_{M}\}$ are configuration parameters determining the increase or decrease of the margin of contribution over time, modeling the progress of the firm in walking the skill learning curve, or the firm’s ability to improve the cost effectiveness of skill development and operations, or the evolution of the market price for the task production under consideration ($\delta_{c}>0$ determines an increase, i.e. a unit of quality-adjusted production increases through epochs, while $\delta_{c}<0$ determines a decrease). Also, we introduce the element of error as a shock conditioned on performance, modeling the economic impact of a low performance not only in terms of reduced output, but also in terms of economic penalty triggered beyond a minimum performance threshold $t_{err}\in[0,1]$. This threshold can be interpreted as a quality control threshold, or a level of safety tolerance. In combination with $t_{err}$, we introduce the parameter $c_{err}\in[0,1]$, which defines the cost of error as a fraction of input in units ($c_{err}$). This represents a variety of phenomena, including the cost of the necessary corrections, the cost of repair for eventual damages, a financial penalty, a reputation damage, or the cost of a product recall. By combining these factors, when $(1-\theta_{i,c,d,e})\geq t_{err}$ we compute $err_{i,c,d,e}=(1-\theta_{i,c,d,e})*c_{err}$, or $err_{i,c,d,e}=0$ otherwise. Finally, the bottom-line economic utility is computed as $u_{i,c,d,e}=v_{i,c,d,e}-err_{i,c,d,e}$. The complete process instructed by the model is more formally describe by Algorithm 4, and schematically illustrated in Figure 1. As schematically illustrated in Figure 2, the result of each simulation $s$ is collected through the following sets of values: $\Theta^{s}=\{\theta_{i,c,d,e,k}^{s}\}$, $Y^{s}=\{y_{i,c,d,e,k}^{s}\}$, $V^{s}=\{v_{i,c,d,e,k}^{s}\}$, $ERR^{s}=\{err_{i,c,d,e,k}^{s}\}$, and $U^{s}=\{u_{i,c,d,e,k}^{s}\}$. For notational convenience, we generically refer to these sets as $\Lambda^{s}=\{\lambda_{i,c,d,e,k}^{s}\}$, with $\lambda\in\{\theta,y,v,err,u\}$ and $\Lambda\in\{\Theta,Y,V,ERR,U\}$. Additionally, we define the subsets of values characterized by the same skill policy $c_{0}$ as $\Lambda^{s}_{c_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0},\forall i,e,d,k\}\subseteq\Lambda^{s}$, and the subsets of values characterized by the same skill policy / task difficulty pair $(c_{0},d_{0})$ as $\Lambda^{s}_{c_{0},d_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0}\land d=d_{0},\forall i,e,k\}\subseteq\Lambda^{s}_{c_{0}}$. Next, for each set $\Lambda^{s}_{[c],[d]}$ (this compact notation stands for $\Lambda^{s}$, $\Lambda^{s}_{c}$ and $\Lambda^{s}_{c,d}$), we compute the arithmetic mean as $\mu(\Lambda^{s}_{[c],[d]})$, the standard deviation as $\sigma(\Lambda^{s}_{[c],[d]})$, the range as $\rho(\Lambda^{s}_{[c],[d]})=\max(\Lambda^{s}_{[c],[d]})-\min(\Lambda^{s}_{[c],[d]})$, the interquartile range (IQR) as $IQR(\Lambda^{s}_{[c],[d]})$, and the skewness as $SK(\Lambda^{s}_{[c],[d]})$. The descriptive statistics $\mu$, $\sigma$, $\rho$, $IQR$ and $SK$ are also generically denoted as $\omega$, crystallizing $\omega(\Lambda^{s}_{[c],[d]})$ as a general notational handles for the characterization and analysis of results.
@@
Algorithm 1 Function $randomchoice:(X,n,P(X))\rightarrow choice$
[…]
2:$n$
[…]
Return: choicechoice $\triangleright$ A vector of length $n$
[…]
3:$n$
[…]
Return: samplesample $\triangleright$ A vector of length $n$
1:param_startparam\_start
2:param_endparam\_end
3:cur_stepcur\_step
4:tot_stepstot\_steps
5:if $total\_steps\leq 1$ then
6:  $cur\_param\leftarrow param\_start$
7:else
9:end if
Return: cur_paramcur\_param

### v6

@@ Abstract
When deploying artificial skills, decision-makers often assume that layering human oversight is a safe harbor that mitigates the risks of full automation in high-complexity tasks. This paper formally challenges the economic validity of this widespread assumption, arguing that the true bottom-line economic utility of a human-machine skill policy is dangerously misunderstood and highly contingent on situational and design factors. To investigate this gap, we develop an in-silico exploratory framework for policy analysis based on Monte Carlo simulations to quantify the economic impact of skill policies in the execution of tasks presenting varying levels of complexity across diverse setups. Our results show that in complex scenarios, a human-machine strategy can yield the highest economic utility, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine approach can perform worse than either the machine-exclusive or the human-exclusive policy, actively destroying value under the pressure of costs that are not sufficiently compensated by performance gains. This finding points to a key implication for decision-makers: when the context is complex and critical, simply allocating human and machine skills to a task may be insufficient, and far from being a silver-bullet solution or a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation when surprise is the norm, even when machines become more effective than humans in handling uncertainty. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of skill policy decisions ex ante. This provides a foundation for scaling experiments and study more general policy implications as candidates for future empirical validation.
@@ 1 Introduction
Research shows that technology can successfully automate routine tasks and complement human skills in more complex work (autor2003skill), while increasingly supplementing human judgment in cognitive tasks (autor2024applying), with supporting experimental evidence in (peng2023impact), (noy2023experimental), and (brynjolfsson2025generative). In this context, organizations of all kinds have a clear opportunity: leverage machine skills to enhance their capabilities and increase competitiveness. Nevertheless, true success hinges not simply in the adoption of new technologies, but on their effective use (agarwal2023combining) (autor2024applying). In turn, this depends not only on the performance achieved by individual skills, whether artificial or human, but also on how they are combined, and on how their use converts into net economic value. However, human and machine skills have distinct characteristics and different performance outcomes across various conditions, as well as different cost and margin structures. This disparity makes the the administrative behavior (Simon1947AdministrativeBehavior) required to determine optimal capital and labor allocations very complex, and risky. We call this the skill policy decision problem, whose solution under uncertainty is far from straightforward.
To understand its underlying nuances, one must first consider the modern evolving landscape of machine capabilities versus their traditional limitations. Recent advancements in methods, combined with increasingly affordable computational capabilities, have allowed machine skills to achieve remarkable performance in tasks where automation has historically struggled. Notable examples include image recognition (krizhevsky2012imagenet), natural language processing (vaswani2017attention), game play (silver2017mastering), skin cancer diagnostics (esteva2017dermatologist), recommendation systems (he2017neural) and protein-folding research (jumper2021highly), just to mention some widely recognized examples. Moreover, models of language deployed through deep neural networks (brown2020language) (hoffmann2022training) (team2023gemini) (touvron2023llama) have contributed to amplify expectations regarding the present and future performance of machine skills in the execution of a wide variety of cognitive tasks. Nevertheless, while machines have proved to excel in learning models on complex and vast data, “crush[ing] the barrier of meaning” (rota1985barrier) in tasks presenting high generalization difficulty (chollet2019measure) can still be challenging for automation. Achieving generalization under uncertainty is a non-monolithic ability involving human-like abstraction and analogy-making (mitchell1993analogy), which can be extremely hard for skill programs to replicate algorithmically. This explains why, as technology improves, the set of tasks that can be automated tends to expand (acemoglu2025simple), but not without limits. Beyond them, in highly uncertain or surprising situations, human skills are often preferred, paying a price in terms of lower efficiency and scalability. In this condition, and especially in the context of cognitive tasks (which are harder to compartmentalize), a potential and rather conventional solution is to avoid the skill policy dilemma altogether and combine human and machine skills to get the best of both. But is this always the safe harbor it seems?
To tackle this question, we build a generative model to analyze the economic effects of skill policy decisions across varying levels of generalization difficulty, investigating the sufficiency of plausible micro-foundational economic assumptions in relation to the emergence of the human-machine paradox. To accomplish this, we combine a task-based framework from the theory of Economics with provisions from Computer Science and Statistics into a model suitable to generate synthetic data through Monte Carlo simulations. This in-silico framework, designed for theoretical explanation and policy exploration (epstein2008model) rather than for empirical prediction, employs bottom-up simulation to quantify emergent insights. The framework operates on the assumption that performance is not a deterministic outcome, but an emergent property shaped by complex interactions between skills and tasks under varying levels of complexity. While the technical challenges of generalization are well-documented in computer science, this work’s primary contribution is not to rediscover that insight. Instead, it introduces a formal method to translate the generalization problem into quantifiable measures of economic utility. To our knowledge, no prior model offers an ex-ante evaluation of skill policies by explicitly bridging the technical concept of generalization difficulty with the economic concepts of error costs and utility. Therefore, this framework is proposed as a novel tool for decision-makers, expanding the discussion from skill performance alone to skill policy decisions, and to how the skill policy of choice needs to be deployed to maximize economic utility.
@@ 2 Model
Our work builds upon a task-based model of production (farboodi2021model) characterizing the quality-adjusted output of a firm $i$ at time $t$ as $y_{i,t}=A_{i,t}k_{i,t}^{\alpha}$, where $k_{i,t}^{\alpha}$ represents the units of output of quality $A_{i,t}$ produced by leveraging $k_{i,t}$ units of capital, and $A_{i,t}=g\big((a_{i,t}-a_{i,t}^{opt})^{2}\big)$ is function of the squared distance between $a_{i,t}$ (the production technique of choice for firm $i$ at time $t$) and the optimal technique $a_{i,t}^{opt}=\theta_{t}+\epsilon_{a,i,t}$}, with $\theta_{t}$ and $\epsilon_{a,i,t}$ being the jointly-observed persistent and transitionary components of the optimal technique, and $g$ a strictly decreasing function. This formulation, by means of the introduction of the squared distance between the actual and the optimal production skill, makes the role of technological performance explicit in the determination of production output of automated tasks. Also, this framework complements the treatments illustrated in (autor2025expertise), where a worker is able to execute a task only if the worker’s expertise level is equal or higher than the one required by the task (the hierarchical expertise assumption), and automation deploys below a certain threshold of task complexity (the hierarchical automation assumption), a condition also illustrated in (acemoglu2025simple) as effect of the comparative advantage offered by labor with respect to capital in newer or complex tasks. Interestingly, this task-based model aligns with the idea of generalization difficulty as presented in (chollet2019measure), i.e. the increased complexity introduced by additional levels of uncertainty presenting at execution time with respect to learning time. The compatibility of these seminal models constitutes the foundational ground for our work.
Moving from this referential framework, we conceptualize performance as a stochastic property that emerges from the interaction between task content and skills, conditioned upon local and global generalization difficulty, extending the definition of production technique to include models operated by the human mind. In this context, we focus on cognitive tasks, leaving the opportunity to relax this assumption in further steps. First, we move by abstraction to defining the distance between the actual and optimal skills in terms of loss ($\mathcal{L}$), such that the quality $A_{i,t}$ can be expressed as $A_{i,t}=g\big(\mathcal{L}_{i,t}\big)$, where $\mathcal{L}_{i,t}=f(a_{i,t},a_{i,t}^{opt})$ and $f$ is any appropriate function measuring the absolute gap between the actual and optimal technological skills. By doing this, we open the possibility to consider alternative measures of distance, including divergence or dissimilarity (also, if necessary, multiple measures of loss may be combined). Then, when the range of potential values of $\mathcal{L}_{i,t}$ is known, we can obtain a normalized loss $\mathcal{L}^{\prime}_{i,t}\in[0,1]$ through the normalization process of choice (for example, $\mathcal{L}^{\prime}_{i,t}=[\mathcal{L}_{i,t}-\min(\mathcal{L}_{i,t})]/[\max(\mathcal{L}_{i,t})-\min(\mathcal{L}_{i,t})]$), and a normalized performance $\theta_{i,t}=(1-\mathcal{L}^{\prime}_{i,t})$. This allows to formally transition from the measure of technological distance to the measure of performance, with consequently easier operational deployment. Conceptually, based on this initial framing, our model operates in a simple sequence. First, a skill attempts a task, resulting in a performance ($\theta$) between zero and one. This performance is then translated into a quality-adjusted output ($y$) and then into an economic value ($v$) by applying a skill-specific margin. Critically, poor performance can also trigger a shock, and the associated cost (errerr). The final economic utility ($u$) is simply the resulting value minus any error cost. The following formalisms detail this process.
Given initial conditions at time $t$, our computational pipeline leverages a stochastic process to determine the normalized performance $\theta_{i,c,d,e}$ obtained by the firm $i$ deploying the skill policy $c$ in the execution of a task with difficulty $d$ at time $t+e$, where $e\in[0,E-1]$ is a task-execution epoch. To this end, without sacrificing generality, we implements three skill policies and three levels of task difficulty. The skill policy options are $H$ (human-exclusive, relying on labor allocations only), $M$ (machine-exclusive, i.e. full automation), and $HM$ (combining human and machine skills), while the levels of task difficulty are set to $High$, MedMed and $Low$. The implementation of skill performance is probabilistic, with distributions $P(\theta)_{c,d,e}$ conditioned on $c$, $d$, and $e$. The global generalization difficulty is modeled by different distributions $P(\theta)_{c,d,e}$ for different levels of difficulty $d$, while the effect of local generalization difficulty is inherently defined by each individual probability distribution $P(\theta)_{c,d,e}$.
At each epoch, the performance obtained by the skill policies $H$ or $M$ (denoted as $\theta_{i,c=H,d,e}$ and $\theta_{i,c=M,d,e}$, and more synthetically as $\theta_{H}$ and $\theta_{M}$, respectively) is randomly drawn according to the appropriate probability distribution $P(\theta)_{c,d,e}$. When, instead, the skill policy adopted by the firm is $HM$, the performance computes as $\theta_{HM}=a(\theta_{H},\theta_{M},\gamma_{HM})$, where $a$ is an arbitrary function defining how the individual skills $H$ and $M$ interact, depending also on an augmenting factor $\gamma_{HM}$. Herein, the concept of augmentation evokes amplification (ashby1956introduction), collaboration (4503259) and integration (engelbart1962augmenting), characterizing a human-machine skill policy in ways that go from task decomposition and sub-task delegation to complex collaborative partnerships.
Then, based on the performance $\theta_{i,c,d,e}$, the quality-adjusted production output computes as $y_{i,c,d,e}=g^{\prime}\big(\theta_{i,c,d,e}\big)$, where $g^{\prime}:\theta_{i,c,d,e}\in[0,1]\rightarrow y_{i,c,d,e}\in[0,1]$ is a monotone and non decreasing function (the combined evaluation of $A_{i,t}$ and ${kl}^{\alpha}_{i,t}$ through the function $g^{\prime}$ is a desired behavior of our model, to streamline the formal relationship between task-execution performance and quality-adjusted output). Next, the value of the production output is determined as $v_{i,c,d,e}=max(0,min(1,y_{i,c,d,e}*mc_{c}*\delta_{c,e}))$, where $mc_{c}\in[0,1],\forall c$ is the margin of contribution associated to each skill policy, and $\delta_{c,e}$, which depends from both $c$ and $e$, represents the skill-specific evolution of the margin of contribution over epochs. To this end, $\delta_{c,e}$ is computed by linear interpolation as $\delta_{c,e}=interpolate(1,1+\delta_{c},e,E)$, according to Algorithm 4, where $\delta_{c}=\{\delta_{H},\delta_{HM},\delta_{M}\}$ are configuration parameters determining the increase or decrease of the margin of contribution over time.
The $\delta_{c}$ parameters describe the progress of the firm in walking the learning curve of each skill, as well as the firm’s ability to improve the cost effectiveness of skill development and operations, or the evolution of the market price for the task production under consideration ($\delta_{c}>0$ determines an increase, i.e. a unit of quality-adjusted production increases through epochs, while $\delta_{c}<0$ determines a decrease). Also, we introduce the element of error as a shock conditioned on performance, modeling the economic impact of a low performance not only in terms of reduced output, but also in terms of economic penalty triggered beyond a minimum performance threshold $t_{err}\in[0,1]$. This threshold can be interpreted as a quality control threshold, or a level of safety tolerance. In combination with $t_{err}$, we introduce the parameter $c_{err}\in[0,1]$, which defines the cost of error as a fraction of input in units ($c_{err}$). This represents a variety of phenomena, including the cost of the necessary corrections, the cost of repair for eventual damages, a financial penalty, a reputation damage, or the cost of a product recall. By combining these factors, when $(1-\theta_{i,c,d,e})\geq t_{err}$ we compute $err_{i,c,d,e}=(1-\theta_{i,c,d,e})*c_{err}$, or $err_{i,c,d,e}=0$ otherwise. Finally, the bottom-line economic utility is computed as $u_{i,c,d,e}=v_{i,c,d,e}-err_{i,c,d,e}$.
@@
The complete process instructed by the model is more formally describe by Algorithm 1, and schematically illustrated in Figure 1. As illustrated in Figure 2, the result of each simulation $s$ is collected through the following sets of values: $\Theta^{s}=\{\theta_{i,c,d,e,k}^{s}\}$, $Y^{s}=\{y_{i,c,d,e,k}^{s}\}$, $V^{s}=\{v_{i,c,d,e,k}^{s}\}$, $ERR^{s}=\{err_{i,c,d,e,k}^{s}\}$, and $U^{s}=\{u_{i,c,d,e,k}^{s}\}$. For notational convenience, we generically refer to these sets as $\Lambda^{s}=\{\lambda_{i,c,d,e,k}^{s}\}$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$ and $\lambda\in\{\theta,y,v,err,u\}$. Additionally, we define the subsets of values characterized by the same skill policy $c_{0}$ as $\Lambda^{s}_{c_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0},\forall i,e,d,k\}\subseteq\Lambda^{s}$, and the subsets of values characterized by the same skill policy / task difficulty pair $(c_{0},d_{0})$ as $\Lambda^{s}_{c_{0},d_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0}\land d=d_{0},\forall i,e,k\}\subseteq\Lambda^{s}_{c_{0}}$. Next, for each set $\Lambda^{s}_{[c],[d]}$ (this compact notation stands for $\Lambda^{s}$, $\Lambda^{s}_{c}$ and $\Lambda^{s}_{c,d}$), we compute the arithmetic mean as $\mu(\Lambda^{s}_{[c],[d]})$, the standard deviation as $\sigma(\Lambda^{s}_{[c],[d]})$, the range as $\rho(\Lambda^{s}_{[c],[d]})=\max(\Lambda^{s}_{[c],[d]})-\min(\Lambda^{s}_{[c],[d]})$, the interquartile range (IQR) as $IQR(\Lambda^{s}_{[c],[d]})$, and the skewness as $SK(\Lambda^{s}_{[c],[d]})$. The descriptive statistics $\mu$, $\sigma$, $\rho$, $IQR$ and $SK$ are also generically denoted as $\omega$, crystallizing $\omega(\Lambda^{s}_{[c],[d]})$ as a general notational handle for the characterization and analysis of results.